OpenStudy (anonymous):

Solve 2x + 5y = 10 for y.

OpenStudy (angela210793):

5y=10-2x y=2-2x/5

OpenStudy (anonymous):

y=(10-2x)/5

OpenStudy (anonymous):

\[y=2-(2/5)x\]